OpenStudy (anonymous):

write an equation parallel to y=3x+5 with x-intercept 10

OpenStudy (anonymous):

Remember that in order for lines to be parallel...they need to have the same slope

OpenStudy (anonymous):

I think the answer is Convert 3x + 2y = 5 into slope-intercept form.. 2y = -3x + 5 y = -3x/2 + 5/2 Since parallel lines contain same slopes, m1 = m2 = -3/2. Hence.. y = -3x/2 - 1 I hope this helps!